Miss Sun Tzi obtained a level principal and interest on the balance loan of $7,265,989. The nominal rate of interest is 0.9% with a term of 5 years and payments scheduled at the end of each day. Compute the interest expense associated with the 16th period.
Answer #1

Nominal interest rate = 0.9%.

Daily interest rate = 0.9%/365.

Number of payments = number of years * 365 = 5 * 365 = 1825.

Interest expense of 16th period is calculated using IPMT function in Excel :

rate = 0.9%/365. (daily interest rate)

per = 16 (we are calculating the interest payment of the 16th period)

nper = 1825 (total number of payments over loan period)

pv = 7265989 (loan amount)

IMPT is calculated to be $177.72

The interest expense associated with the 16th period is $177.72.

A11 fx =IPMT(0.9%/365,16,1825,7265989) D E F G B C 11 (177.72)